(a)    Suppose widgets have a price elasticiy of demand equal to -1.95 and a cross elasticity of demand equal to 2 with dundles. What are the pricing options for a firm that sells widgets? Explain.

(b)   If the quantity demand for apples is given by:

                                                    QD = 300 − 2.55pA + 7.62pB + 0.75Y            

if pA = 90 cents, pB = 50cents and QD = 860.

Calculate and interpret the income elasticity of demand for apples.


1
Expert's answer
2021-07-27T03:58:01-0400

(a)

Firms that sell widgets should set the price of widgets to be lower than that of dundles.

This is because the two goods are substitutes and widgets are relatively elastic in this case.

(b)

"Q_D=300-2.55P_A+7.62P_B+0.75Y"

"860=300-2.55(90)+7.62(50)+0.75Y"

"408.5=0.75Y"

"Y=544.67"

Income elasticity:

="\frac {544.67}{860}"

=0.6333

This implies that apples are normal goods because their income elasticity falls between 0 and 1.Demand for apples is an income inelastic demand.